OpenStudy (anonymous):

Find the product of (3x + 7y)2.

OpenStudy (anonymous):

any options?

OpenStudy (funnyguy2):

(3x+7y)(2) =(3x+7y)(2) =(3x)(2)+(7y)(2) =6x+14y

OpenStudy (anonymous):

yes gimme some time to post them its a lot

OpenStudy (anonymous):

that's not the right answer

OpenStudy (anonymous):

a.)9x^2 +42xy +49y^2

OpenStudy (anonymous):

b.)9x^2 - 49y^2

OpenStudy (anonymous):

c.)9x^2 +49y^2

OpenStudy (anonymous):

d.)9x^2 +42xy 49y^2

OpenStudy (mathstudent55):

The 2 is an exponent, right?

OpenStudy (anonymous):

the is is squared

OpenStudy (anonymous):

2

OpenStudy (mathstudent55):

\((3x + 7y)^2\) \(= (3x + 7y)(3x + 7y)\) Use FOIL

OpenStudy (anonymous):

? its multiple choice I posted abc or d

OpenStudy (mathstudent55):